# Weak convergence of Bayes estimators under general loss functions

### v1

@@ C.2 Proofs for Section 3
for any $t\in K$. Thus for any $t\in K$ and any $h\in H_{n}$ there exists some $\xi>0,0<a<2$ such that
@@
This finishes the proof. ∎
@@ C.3 Technical details for Example 1
Set
$0<\lambda_{\text{min}}:=\inf_{\vartheta\in B_{\varepsilon}(\vartheta_{0}),u\in S^{d-1}}D^{k}(\vartheta)[u]\leq\lambda_{\text{max}}:=\sup_{\vartheta\in B_{\varepsilon}(\vartheta_{0}),u\in S^{d-1}}D^{k}(\vartheta)[u]<\infty,$
where $S^{d-1}=\{u\in\mathbb{R}^{d}:\left\lVert u\right\rVert=1\}$. As $D^{k}(\vartheta)[u]$ is homogeneous in $u\in\mathbb{R}^{d}$ of order $k$, we have
@@
for any $\vartheta\in B_{\varepsilon}(\vartheta_{0})$. By a Taylor expansion of order $k$
@@
Choosing $\varepsilon<\lambda_{\text{min}}/(2M)$ gives the lower bound. The upper bound is obtained for $\varepsilon<1$ by
@@
C.4 Proofs for Section 4
Proof of Theorem 3.
We want to apply Theorem 2 for $\ell_{0}=\left\langle t-h,\,A(\vartheta_{0})(t-h)\right\rangle$, which would yield the result. The only conditions left to show is condition (iii) from Theorem 2. For this purpose, we want to use Proposition 3. Condition (i), (ii) and (iv) from Proposition 3 is satisfied, because of Assumptions 3 and 4. The only thing left to show is that
$\sup_{t\in K}\left\lvert\frac{\ell(t\varepsilon+\vartheta_{0},h\varepsilon+\vartheta_{0})}{\varepsilon^{2}}-\ell_{0}(t,h)\right\rvert\to 0,\quad\varepsilon\to 0$
for any $h\in\mathbb{R}^{d}$. For any $\delta>0$ there exists $\varepsilon>0$ small enough such that $t\varepsilon+\vartheta_{0},h\varepsilon+\vartheta_{0}\in B_{\delta}(\vartheta_{0})$ for all $t\in K$. Choosing $\delta>0$ small enough such that $\left\lvert\xi(t,\vartheta)\right\rvert\lesssim\left\lVert t-\vartheta\right\rVert^{c_{1}}$ for some $c_{1}>2$ and any $t,\vartheta\in B_{\delta}(\vartheta_{0})$ yields
@@ C.5 Technical derivations for Section 5
Additional details in Example 4. We show the claim in Example 4: For any compact $K\subset\mathbb{R}^{d}$ and any $h\in\mathbb{R}^{d}$
@@
For ease of notation let $t_{\varepsilon}=t\varepsilon+\vartheta_{0},h_{\varepsilon}=h\varepsilon+\vartheta_{0}$ and $g_{\vartheta}(t)=\left\lVert p_{t}-p_{\vartheta}-\left\langle t-\vartheta,\,\dot{p}_{\vartheta}\right\rangle\right\rVert_{L^{1}}$. Note that $L^{1}$ differentiability implies that $g_{h}(t)=o(\left\lVert t-h\right\rVert)$, as $t\to h$. Thus, for any $t\in K$, where $K\subset\mathbb{R}$ is any compact set it holds
@@
$\sup_{t\in K}\frac{g_{\vartheta_{0}}(t_{\varepsilon})}{\varepsilon}=\sup_{t\in K}\varepsilon^{-1}\left\lVert p_{t_{\varepsilon}}-p_{\vartheta_{0}}-\left\langle t_{\varepsilon}-\vartheta_{0},\,\dot{p}_{\vartheta_{0}}\right\rangle\right\rVert_{L^{1}}\to 0.$
For an arbitrary, but fixed $x\in\mathcal{X}$ denote
@@
Note that $f_{\varepsilon}(t)$ converges to zero for any $t$, by differentiability of the normal density in the parameter. If we can show that it also holds uniformly in $t$, we conclude that also $\sup_{t\in K}\varepsilon^{-1}g_{\vartheta_{0}}(t_{\varepsilon})\to 0$, by the dominated convergence theorem. The required property is directional differentiability of $\vartheta\mapsto p_{\vartheta}(x)$ at $\vartheta_{0}$ in direction $t$ uniformly over all $t\in K$. This follows from the fact that the normal density is continuously differentiable in the parameter: For $\eta=\sup_{t\in K}\left\lVert t\right\rVert<\infty$ and for some $0\leq\lambda\leq 1$ the mean value theorem yields
$f_{\varepsilon}(t)=\left\lvert\left\langle t,\,\dot{p}_{\vartheta_{0}+\eta\lambda t}(x)-\dot{p}_{\vartheta_{0}}(x)\right\rangle\right\rvert\leq\eta\sup_{\vartheta\in B_{\eta\varepsilon}(\vartheta_{0})}\left\lVert\dot{p}_{\vartheta}(x)-\dot{p}_{\vartheta_{0}}(x)\right\rVert.$
The upper bound converges to zero, as $\varepsilon\to 0$ by continuity.
Additional details in Section 5.3.
@@ C.6 Technical derivations for Section 6
Denote ${I}_{d}$ as the $d\times d$ identity matrix. Then, we have for any feasible $\pi\in[0,1]^{d\times d}$ that

### v2

@@ C.2 Proofs for Section 3
for any $t\in K$. Thus, for any $t\in K$ and any $h\in H_{n}$ there exist some $\xi>0,0<a<2$ such that
@@
Now the assertion follows directly from Lemma 1. ∎
@@ C.3 Technical details for Example 1
Let $S^{d-1}=\{u\in\mathbb{R}^{d}:\left\lVert u\right\rVert=1\}$ and set
$0<\lambda_{\text{min}}:=\inf_{\vartheta\in B_{\varepsilon}(\vartheta_{0}),u\in S^{d-1}}D^{k}(\vartheta)[u]\leq\lambda_{\text{max}}:=\sup_{\vartheta\in B_{\varepsilon}(\vartheta_{0}),u\in S^{d-1}}D^{k}(\vartheta)[u]<\infty.$
As $D^{k}(\vartheta)[u]$ is homogeneous in $u\in\mathbb{R}^{d}$ of order $k$, we have
@@
for any $\vartheta\in B_{\varepsilon}(\vartheta_{0})$. By a Taylor expansion of order $k$,
@@
Choosing $\varepsilon<\lambda_{\text{min}}/(2M)$ yields the lower bound. The upper bound is obtained, for $\varepsilon<1$, by
@@
C.4 Technical details for Section 4
C.4.1 Detailed derivation for equation (5)
The upper bound follows from
For the lower bound, let $\lambda_{\min}(\vartheta)$ denote the smallest eigenvalue of $A(\vartheta)$. By continuity of $\vartheta\mapsto A(\vartheta)$ and positive definiteness of $A(\vartheta_{0})$, there exists $\delta>0$ such that
and $\eta_{\delta}$ is nonincreasing in $\delta$, we can choose $\varepsilon>0$ and $\eta>0$ such that $\eta_{\delta}\geq\eta>0$ for all $\delta<\varepsilon$. If $\delta<\varepsilon$ and $t,\vartheta\in B_{\delta}(\vartheta_{0})$, then
C.4.2 Proof of Theorem 3
We apply Theorem 2 with $\ell_{0}=\left\langle t-h,\,A(\vartheta_{0})(t-h)\right\rangle$, which yields the result. It therefore suffices to verify requirement (iii) in Theorem 2.
To this end, we invoke Proposition 3, for which requirement (i) follows from Assumption 3, particularly, from equation (5), whose derivation is given above. Requirements (ii) and (iv) in Proposition 3 are satisfied by Assumptions 3 and 4. It remains to verify requirement (iii) in Proposition 3, namely, for any $h\in\mathbb{R}^{d}$,
$\sup_{t\in K}\left\lvert\frac{\ell(t\varepsilon+\vartheta_{0},h\varepsilon+\vartheta_{0})}{\varepsilon^{2}}-\ell_{0}(t,h)\right\rvert\to 0,\quad\varepsilon\to 0.$
For any $\delta>0$ there exists $\varepsilon>0$ suffiiciently small such that $t\varepsilon+\vartheta_{0},h\varepsilon+\vartheta_{0}\in B_{\delta}(\vartheta_{0})$ for all $t\in K$. Choose $\delta>0$ small enough and some $c_{1}>2$, such that $\left\lvert\xi(t,\vartheta)\right\rvert\lesssim\left\lVert t-\vartheta\right\rVert^{c_{1}}$ for all $t,\vartheta\in B_{\delta}(\vartheta_{0})$. Then
@@ C.5 Technical derivations for Section 5
C.5.1 Additional details in Example 3
The reverse triangle inequality and the Cauchy–Schwarz inequality yield uniform convergence as follows
which converges to zero, as $\varepsilon\to 0$, by continuity.
C.5.2 Additional details in Example 4
We show the claim in Example 4: For any compact $K\subset\mathbb{R}^{d}$ and any $h\in\mathbb{R}^{d}$
@@ C.5.2 Additional details in Example 4
For ease of notation, let $t_{\varepsilon}=t\varepsilon+\vartheta_{0},h_{\varepsilon}=h\varepsilon+\vartheta_{0}$ and $g_{\vartheta}(t)=\left\lVert p_{t}-p_{\vartheta}-\left\langle t-\vartheta,\,\dot{p}_{\vartheta}\right\rangle\right\rVert_{L^{1}}$. Note that $L^{1}$-differentiability implies that $g_{h}(t)=o(\left\lVert t-h\right\rVert)$, as $t\to h$. Thus, for any $t\in K$, where $K\subset\mathbb{R}$ is any compact set, it holds
@@
$\sup_{t\in K}\frac{g_{\vartheta_{0}}(t_{\varepsilon})}{\varepsilon}=\sup_{t\in K}\frac{\left\lVert p_{t_{\varepsilon}}-p_{\vartheta_{0}}-\left\langle t_{\varepsilon}-\vartheta_{0},\,\dot{p}_{\vartheta_{0}}\right\rangle\right\rVert_{L^{1}}}{\varepsilon}\to 0.$
For an arbitrarily fixed $x\in\mathcal{X}$, denote
@@
Note that $f_{\varepsilon}(t)$ converges to zero for any $t$, by differentiability of the normal density in the parameter. If we can show that it also holds uniformly in $t$, we conclude that also $\sup_{t\in K}\varepsilon^{-1}g_{\vartheta_{0}}(t_{\varepsilon})\to 0$, by the dominated convergence theorem. The required property is directional differentiability of $\vartheta\mapsto p_{\vartheta}(x)$ at $\vartheta_{0}$ in direction $t$ uniformly over all $t\in K$. This follows from the fact that the normal density is continuously differentiable in the parameter: For $\eta=\sup_{t\in K}\left\lVert t\right\rVert<\infty$ and $0\leq\alpha\leq 1$, the mean value theorem yields
$f_{\varepsilon}(t)=\left\lvert\left\langle t,\,\dot{p}_{\vartheta_{0}+\eta\alpha t}(x)-\dot{p}_{\vartheta_{0}}(x)\right\rangle\right\rvert\leq\eta\sup_{\vartheta\in B_{\eta\varepsilon}(\vartheta_{0})}\left\lVert\dot{p}_{\vartheta}(x)-\dot{p}_{\vartheta_{0}}(x)\right\rVert.$
The upper bound converges to zero, as $\varepsilon\to 0$, by continuity.
C.5.3 Additional details in Section 5.3
Note that
@@ C.6 Technical derivations for Section 6
Let ${I}_{d}$ be the $d\times d$ identity matrix. Then, we have, for any feasible $\pi\in[0,1]^{d\times d}$,